OpenStudy (anonymous):

What is the eighth term of the geometric sequence 3, 12, 48, 192, ... ?

OpenStudy (amoodarya):

a1=3 a2=12=a1*q so q=4 then a(n)=a1 q^(n-1) a(8)=a1*q^7=3*(4)^7

OpenStudy (amoodarya):

a7=3*1638=49152